OpenStudy (anonymous):

is this an identity? 10tanX/1+tan^2X=5sinX i don't think it is.....?

OpenStudy (ash2326):

Nope. not a standard identity

OpenStudy (anonymous):

i tried alot of different things but i thought i might have done something wrong

OpenStudy (anonymous):

\[\frac{10\tan ^{2} x}{\sec ^{2}x}\]change to sines and cosines and reduce

OpenStudy (ash2326):

in fact it's wrong \[\frac{10\tan x}{1+\tan^2 x}\] \[\frac{10\tan x}{1+\frac{\sin ^2 x}{\cos^2 x}}\] \[\frac{10\tan x\times \cos^2 x}{\cos^2 x+\sin ^2 x}\] \[\frac{10\tan x\times \cos^2 x}{1}\] \[10\times \frac{\sin x}{\cos x}\times \cos^2 x\] \[10\times \sin x\cos x\] \[5\times \underline {2 \sin x \cos x}\] \[5\times \sin 2x\]

OpenStudy (ash2326):

@jotwestmoreland do you get this?

OpenStudy (anonymous):

let x=pi/4... left side is \(\large \frac{10tan(\pi /4)}{1+tan^2(\pi /4)}=5 \) right side is \(\large 5 \cdot sin(\pi /4)=\frac{5\sqrt2}{2} \)

OpenStudy (anonymous):

There you go... lot's of ways to show that it just won't work.

OpenStudy (anonymous):

if it was an identity, the left = the right ....

OpenStudy (anonymous):

*for any x...

OpenStudy (anonymous):

thank you all so much for confirming that!!

OpenStudy (ash2326):

:D welcome